Build a rectangular pen with three parallel partitions using 500 feet of fencing. What dimensions will maximize the total area of the pen 


1
Expert's answer
2021-11-14T18:17:12-0500

Diagram is missing.

Adding diagram of the problem:



Solution:

Let , width be "x" and length be "y" .

The total amount of fencing is given by

"500=5x+2y\\
2y=500-5x\\
y=250-(5/2)x"

Maximum area "(A)=xy"

"A=x(250-(5/2)x)\\
A=250x-(5/2)x^2"

Diff. w.r.t "x"

"\frac{dA}{dx}=\frac{d}{dx}(250x- \frac{5}{2}x^2)
=250-5x"

For maximum area:

"250-5x=0\Rightarrow x=50"

"y=250-(5\/2)x=250-(5\/2)(50)=125"

Hence, width is 50 ft and length is 125 ft.